OpenStudy (anonymous):

Solve the equation. log(x+5)-logx=3 please help me

OpenStudy (anonymous):

Log ( a) - log ( b ) = log ( a/b )

OpenStudy (anonymous):

oh gosh.. um log 5/3?

OpenStudy (anonymous):

Log ( (x+ 3)/ x) = 3

OpenStudy (anonymous):

(x + 3)/x = 10^3